# High-dimensional limit theorems for SGD: Momentum and Adaptive Step-sizes

### v1

@@ 6.4. Dynamics for SGD-U (Even Powered Monomials)
Let $0<\gamma\leq 0.2$. By Lemma 6.3, 6.4, and the extension of $c(0):=0$, $c$ is continuous on the compact interval $0,\gamma]$, hence attains a maximum $C_{\gamma}=\max_{[0,\gamma]}c(r)$.
@@ 7.1. Norms of Gaussian Vectors
Noting that $\|X\|^{2}=X^{\top}X$ is a quadratic form, the cumulants can be determined exactly, for example see [magnus1986exact, Lemma 2]:
@@
As $\|x\|$ is a radial function, by the coarea formula (e.g. see [stromberg2015introduction, pg. 370]) the remaining integral becomes
@@
Let $Z\sim\mathcal{N}(0,I_{n})$ be a standard Gaussian vector. We first determine a comparison inequality involving $\mathbb{E}\|X\|^{-2k}$ for $\mu=0$, and show that the upper bound is still applicable $\mu\neq 0$ by Anderson’s inequality [anderson1955integral]. In particular, we have that for the centered Gaussian density $\gamma_{\Sigma}$ and a origin-symmetric, convex set $K\subset\mathbb{R}^{n}$
@@
by layer-cake representation. Note that the inner integral is over the centered Euclidean ball
@@
It remains to determine the asymptotic rate for $\mathbb{E}[\|Z\|^{-2k}]$. Since $\|Z\|^{2}$ follows a $\chi^{2}_{n}$-distribution and $n>2k$ with $k\geq 1$, the inverse moments are given by [soch2023statproofbook]

### v2

@@ 8.2. Fixed Point Analysis for SGD-M (Monomial)
8.3. Diffusive Limits for SGD-M
We present a more general version of Proposition 3.12 below.
Theorem 8.1.
Suppose the link function $f\in C^{2}$ satisfies $\mathbb{E}_{Z}[f^{\prime}(Z)]=0$ for $Z\sim\mathcal{N}(0,1)$. Then for SGD-M, the rescaled statistics $\tilde{\boldsymbol{u}}_{n}(t)=(\tilde{m},r^{2})=(\sqrt{n}m,r^{2})$ converges as $n\to\infty$ to the solution of the following SDE,
$\text{d}\tilde{m}$ $=-\frac{2}{1-\beta}\tilde{m}\mathbb{E}\left[a_{1}^{2}\left(f^{\prime\prime}(ra_{2})(f(ra_{2})-f(a_{1}))+f^{\prime}(ra_{2})^{2}\right)\right]\,\text{d}t$
$\qquad\qquad+\frac{1}{1-\beta}\sqrt{4c_{\delta}\mathbb{E}\left[a_{1}^{2}f^{\prime}(ra_{2})^{2}\left((f(ra_{2})-f(a_{1}))^{2}+\sigma^{2}\right)\right]}\,\text{d}B_{t}$
$\text{d}r^{2}$ $=-\frac{4}{1-\beta}\bigg[r\mathbb{E}\left[a_{2}f^{\prime}(ra_{2})(f(ra_{2})-f(a_{1}))\right].$
(21) $\qquad\qquad-\frac{c_{\delta}}{1-\beta}\mathbb{E}\left[f^{\prime}(ra_{2})^{2}\left((f(ra_{2})-f(a_{1}))^{2}+\sigma^{2}\right)\right]\big]\,\text{d}t$
where $a_{1},a_{2}\sim\mathcal{N}(0,1)$ independently and $B_{t}$ is a standard one-dimensional Brownian motion.
Proof.
From Proposition 3.5, we rewrite the drift components for the dynamics associated with $\tilde{u}_{2}=r^{2}$ as
$\langle\nabla\Phi,\nabla r^{2}\rangle$ $=-4r\,\mathbb{E}\left[a_{2}f^{\prime}\left(a_{1}(\tilde{u}_{1}/\sqrt{n})+a_{2}r\right)\left(f\left(a_{1}(\tilde{u}_{1}/\sqrt{n})+a_{2}r\right)-f(a_{1})\right)\right]$
$\frac{c_{\delta}}{2n}\langle V,\nabla^{2}r^{2}\rangle$ $=4c_{\delta}\,\mathbb{E}\left[f^{\prime}\left(a_{1}(\tilde{u}_{1}/\sqrt{n})+a_{2}r\right)^{2}\left((f\left(a_{1}(\tilde{u}_{1}/\sqrt{n})+a_{2}r\right)-f(a_{1}))^{2}+\sigma^{2}\right)\right]$
Taking the limit as $n\to\infty$ recovers the drift components for $r^{2}$. For the dynamics of $\tilde{u}_{1}=\sqrt{n}m$, we linearize the drift about $m=0$,
$F_{m}(m,r)$ $=\mathbb{E}[a_{1}f^{\prime}(a_{1}m+a_{2}r)(f(a_{1}m+a_{2}r)-f(a_{1}))]$
$=\mathbb{E}[a_{1}f^{\prime}(a_{2}r)(f(a_{2}r)-f(a_{1}))]+m\partial_{m}F(0,r)+o(m)$
By independence of $a_{1},a_{2}$, the leading term reduces to
$\mathbb{E}[a_{1}f^{\prime}(a_{2}r)f(a_{2}r)]-\mathbb{E}[a_{1}f^{\prime}(a_{2}r)f(a_{1})]=-\mathbb{E}[a_{1}f(a_{1})]\mathbb{E}[f^{\prime}(a_{2}r)]$
By Stein’s lemma, we have that $\mathbb{E}[a_{1}f(a_{1})]=\mathbb{E}[f^{\prime}(a_{1})]$ which is zero by assumption. For the first-order term, the partial with respect to $m$ is given by
$\partial_{m}F(m,r)=\mathbb{E}[a_{1}^{2}f^{\prime\prime}(a_{1}m+a_{2}r)(f(a_{1}m+a_{2}r)-f(a_{1}))+f^{\prime}(a_{1}m+a_{2}r)^{2}]$
Thus under the rescaling of $\tilde{u}_{1}=\sqrt{n}m$, we have that
$\langle\nabla\Phi,\nabla\tilde{u}_{1}\rangle=\sqrt{n}\left[\frac{\tilde{u}_{1}}{\sqrt{n}}\mathbb{E}[a_{1}^{2}(f^{\prime\prime}(a_{2}r)(f(a_{2}r)-f(a_{1}))+f^{\prime}(a_{2}r)^{2})]+o(n^{-1/2})\right]$
which recovers the limiting drift for $\tilde{u}_{1}$ as $n\to\infty$. For the diffusion term, recall the limiting volatility matrix from the proof of Proposition 3.5,
where we set $s:=\langle x,g\rangle=ma_{1}+ra_{2}$ and $\Delta:=f(s)-f(a_{1})+\varepsilon$. Evidently, under the rescaling only $\Sigma_{11}$ survives the $n\to\infty$ limit as reduces to
$\Sigma_{11}(\tilde{u}_{1},r)=4\mathrm{Var}\left[\left(f(ra_{2})-f(a_{1})+\varepsilon\right)f^{\prime}(ra_{2})a_{1}\right]$
To simplify the variance, note that
$\mathbb{E}\left[\left(f(ra_{2})-f(a_{1})+\varepsilon\right)f^{\prime}(ra_{2})a_{1}\right]$ $=-\mathbb{E}[a_{1}f(a_{1})]\mathbb{E}[f^{\prime}(ra_{2})]=0$
where the reduction follows by independence of $a_{1},a_{2},\epsilon$ followed by our assumption that $\mathbb{E}[a_{1}f(a_{1})]=\mathbb{E}[f^{\prime}(a_{1})]=0$ by Stein’s lemma. Thus the limiting volatility coefficient as $n\to\infty$ is
$\Sigma_{11}(\tilde{u}_{1},r)=4\mathbb{E}\left[a_{1}^{2}f^{\prime}(ra_{2})^{2}\left((f(ra_{2})-f(a_{1}))^{2}+\sigma^{2}\right)\right]$
which completes the proof. ∎
The condition $\mathbb{E}_{Z}[f^{\prime}(Z)]=0$ is to ensure that the fixed point for which we are considering the diffusive limit exists along $m=0$. This holds, for example, if $f$ is an even function such as in phase retrieval where $f(x)=x^{2}$.
Proof of Proposition 3.12.
With Theorem 8.1, we compute each of the coefficients using $f(x)=x^{2}$, $f^{\prime}(x)=2x$, and $f^{\prime\prime}(x)=2$,
$2\mathbb{E}\left[a_{1}^{2}\left(f^{\prime\prime}(ra_{2})(f(ra_{2})-f(a_{1}))+f^{\prime}(ra_{2})^{2}\right)\right]$ $=2\mathbb{E}[a_{1}^{2}(6r^{2}a_{2}^{2}-2a_{1}^{2})]=12(r^{2}-1)$
$4r\mathbb{E}\left[a_{2}f^{\prime}(ra_{2})(f(ra_{2})-f(a_{1}))\right]$ $=8r^{2}\mathbb{E}[a_{2}^{2}(r^{2}a_{2}^{2}-a_{1}^{2})]=8r^{2}(3r^{2}-1)$
and with $\sigma\to 0$,
$4\mathbb{E}\left[f^{\prime}(ra_{2})^{2}\left((f(ra_{2})-f(a_{1}))^{2}+\sigma^{2}\right)\right]$ $=16r^{2}\mathbb{E}[a_{2}^{2}(r^{2}a_{2}^{2}-a_{1}^{2})^{2}]+16r^{2}\sigma^{2}$
$=16r^{2}(15r^{4}-6r^{2}+3+\sigma^{2})$
$4\mathbb{E}\left[a_{1}^{2}f^{\prime}(ra_{2})^{2}\left((f(ra_{2})-f(a_{1}))^{2}+\sigma^{2}\right)\right]$ $=16r^{2}\mathbb{E}[a_{1}^{2}a_{2}^{2}(r^{2}a_{2}^{2}-a_{1}^{2})^{2}]+16r^{2}\sigma^{2}$
$=16r^{2}(15r^{4}-18r^{2}+15+\sigma^{2})$
Substituting the quantities recovers the diffusive dynamics in (3.12). ∎
@@ 8.5. Dynamics for SGD-U (Even Powered Monomials)
Let $0<\gamma\leq 0.2$. By Lemma 8.6, 8.7, and the extension of $c(0):=0$, $c$ is continuous on the compact interval $0,\gamma]$, hence attains a maximum $C_{\gamma}=\max_{[0,\gamma]}c(r)$.
@@ 9.1. Norms of Gaussian Vectors
Noting that $\|X\|^{2}=X^{\top}X$ is a quadratic form, the cumulants can be determined exactly, for example see [30, Lemma 2]:
@@
As $\|x\|$ is a radial function, by the coarea formula (e.g. see [50, pg. 370]) the remaining integral becomes
@@
Let $Z\sim\mathcal{N}(0,I_{n})$ be a standard Gaussian vector. We first determine a comparison inequality involving $\mathbb{E}\|X\|^{-2k}$ for $\mu=0$, and show that the upper bound is still applicable $\mu\neq 0$ by Anderson’s inequality [1]. In particular, we have that for the centered Gaussian density $\gamma_{\Sigma}$ and a origin-symmetric, convex set $K\subset\mathbb{R}^{n}$
@@
by “layer-cake” representation. Note that the inner integral is over the centered Euclidean ball
@@
It remains to determine the asymptotic rate for $\mathbb{E}[\|Z\|^{-2k}]$. Since $\|Z\|^{2}$ follows a $\chi^{2}_{n}$-distribution and $n>2k$ with $k\geq 1$, the inverse moments are given by [49]